Problem 1

Below, you are provided with the demand and supply curves for t-shirts and the world price of a t-shirt. You will usethis information to identify whether the country imports or exports t-shirts. You will also examine the impact of a tariffon the amount of consumer and producer surplus that results in this market.

Price .- .--.-- -- --- --- Supply - - - - - World Price Demand 100 200 300 400 500 600 Quantity of V-Neck T-Shirts

  1. Suppose that the world price of a t-shirt is $20. Does this country import or export t-shirts? How many?
  2. Suppose that this country engages in trade.
  1. Calculate the amount of consumer surplus that results in this market.
  2. Calculate the amount of producer surplus that results in this market.
  1. Suppose that the government imposes a $10 import tariff on t-shirts.
  1. How many t-shirts do domestic consumers purchase after this import tariff is imposed?
  2. How many t-shirts do domestic suppliers sell after this import tariff is imposed?
  3. How many t-shirts does this county import?
  4. Calculate the amount of consumer surplus that results in this market.
  5. Calculate the amount of producer surplus that results in this market.
  6. Calculate the amount of government tax revenue generated by the import tariff.
  7. Calculate the amount of deadweight loss that results from the imposition of the tariff.
Answer #1

Price Supply: Consumer Surplus.(cs World Price Progucer Surplu Imports Demand 100 200 300 400 500 600 Quantity of V-Neck T-Sh

At price of $ 20, Imports = Quantity demanded - Quantity supplied = 400 - 100 = 300 T-shirts

a) Consumer Surplus = 1/2 x base x height = 1/2 x 400 x (60 - 20) = 200 x 40 = 8,000

b) Producer Surplus = 1/2 x 100 x (20 - 0) = 50 x 20 = 1,000

Price after tariff = 20 + 10 = $ 30

Price Supply WP + Tariff Govt. World Price $20 PS bw revenue DWL Demand 100 200 300 400 500 600 Quantity of V-Neck T-Shirts

a) 300 t-shirts are purchased by domestic consumers.

b) 150 t-shirts are supplied by domestic suppliers.

c) Imports = Quantity demanded - Quantity supplied = 300 - 150 = 150 t-shirts

d) Consumer Surplus = 1/2 x 300 x (60 - 30) = 150 x 30 = 4500

e) Producer Surplus = 1/2 x 150 x (30 - 0) = 150 x 15 = 2250

f) Government tax revenue = (30 - 20)(300 - 150) = 10 x 150 = 1500

g) DWL = 1/2 x (30 - 20) x (150 - 100) + 1/2 x (30 - 20) x (400 - 300) = 5 x 50 + 5 x 100 = 250 + 500 = 750